### v1

# Exact Expressive Power of Transformers with Padding

Abstract
Chain of thought is a natural inference-time method for increasing the computational power of transformer-based large language models (LLMs), but comes at the cost of sequential decoding. Are there more efficient alternatives to expand a transformer’s expressive power without adding parameters? We consider transformers with padding tokens as a form of parallelizable test-time compute. We show that averaging-hard-attention, masked-pre-norm transformers with polynomial padding converge to precisely the class $\mathsf{TC}^{0}$ of extremely parallelizable problems. While the $\mathsf{TC}^{0}$ upper bound was known, proving a matching lower bound had been elusive. Further, our novel analysis reveals the precise expanded power of padded transformers when coupled with another form of inference-time compute, namely dynamically increasing depth via looping. Our core technical contribution is to show how padding helps bring the notions of complete problems and reductions, which have been a cornerstone of classical complexity theory, to the formal study of transformers. Armed with this new tool, we prove that padded transformers with $O(\log^{d}n)$ looping on inputs of length $n$ recognize exactly the class $\mathsf{TC}^{d}$ of moderately parallelizable problems. Thus, padding and looping together systematically expand transformers’ expressive power: with polylogarithmic looping, padded transformers converge to the class $\mathsf{NC}$, the best that could be expected without losing parallelism (unless $\mathsf{NC}=\mathsf{P}$). Our results thus motivate further exploration of padding and looping as parallelizable alternatives to chain of thought.
1 Introduction
Due to the computational limitations of transformers (Merrill and Sabharwal, 2023a; Strobl et al., 2024; Chiang, 2025), solving complex reasoning problems requires extending their computational power at inference time, typically by allowing models to generate long chains of thought (CoT) before their outputs (Wei et al., 2022; Nye et al., 2022). Theoretical work has shown how CoT expands the expressive power of transformers to sequential problems outside $\mathsf{TC}^{0}$, but it also sacrifices parallelism (Merrill and Sabharwal, 2023b; Li et al., 2024), making inference slow. Are there alternative inference-time compute approaches that increase expressive power while preserving parallelism?
One method for parallelizable inference-time compute with LLMs is using padding tokens rather than CoT. Padding can be understood as a restriction of CoT where the tokens on the chain are restricted to some “blank” symbol rather than tokens generated by the LLM. Since all the input tokens are known in advance, this makes padding more parallelizable than CoT. There have been some attempts to make padding practical with mixed results (Goyal et al., 2024; Pfau et al., 2024), but it is not fully understood. Specifically, while it is known that padded transformers are in $\mathsf{TC}^{0}$, it has been open and elusive whether they can solve all problems in $\mathsf{TC}^{0}$ or even the smaller class $\mathsf{AC}^{0}$ (Pfau et al., 2024).
Our first contribution in this work is exactly characterizing the expressive power of transformers with polynomial padding as $\mathsf{TC}^{0}$, answering an open question (Pfau et al., 2024). The result emerges through a finer-grained analysis of transformers in terms of string logics (Merrill and Sabharwal, 2023c; Chiang et al., 2023; Yang et al., 2024; Yang and Chiang, 2024). Thinking in terms of logic, we show that $n^{k}$ padding tokens give transformers enough “storage space” to resolve any first-order majority logic formula over $k$ variables. This suffices to capture all of $\mathsf{TC}^{0}$ (Mix Barrington et al., 1990), giving an exact characterization of the expressive power of such transformers.
This first result, however, does not clarify whether padded transformers can be minimally extended to gain expressivity beyond $\mathsf{TC}^{0}$. To address this, we consider the combination of padding with looping, i.e., repeating a block of layers dynamically as a function of input length (Dehghani et al., 2019; Giannou et al., 2023). This can be understood as a form of inference-time compute where padding controls the computation width and looping controls the computation depth—crucially, without adding any parameters. If the number of repetitions is minimal (e.g., sublinear) in sequence length, this transformer model remains highly parallelizable relative to CoT (Merrill and Sabharwal, 2024).
We extend our first result to give an exact characterization of the expressive power of padded transformers with minimal looping. Specifically, transformers with polynomial padding and $O(\log^{d}n)$ looping recognize exactly $\mathsf{TC}^{d}$. This means that log or polylog-looped transformers can solve many problems outside $\mathsf{TC}^{0}$ under standard complexity conjectures, including boolean formula evaluation, iterated matrix multiplication, graph connectivity, and context-free language recognition. It also follows that polylog-looped padded transformers converge in expressive power to $\mathsf{NC}$, the ceiling that could be expected while preserving parallelism under standard complexity conjectures.
Our first result about the expressive power of fixed-depth padded transformers provides crucial insight for our extended results about looped padded transformers (cf. Figure 1), by allowing, for the first time, the use of familiar tools of reductions and complete problems from classical complexity theory in the analysis of transformers. First, the ability to express $\mathsf{TC}^{0}$ implies that padded transformers can implement $\mathsf{FO}$ reductions. We then scaffold the ability to implement $\mathsf{FO}$ reductions to show that transformers can also implement $\mathsf{L}$ reductions, via known results about transformers’ ability to implement graph connectivity, an $\mathsf{NL}$-complete problem under $\mathsf{FO}$ reductions. We then use this to show that $O(\log^{d}n)$ looping recognizes exactly $\mathsf{L}$-uniform $\mathsf{TC}^{d}$, via the $\mathsf{TC}^{d}$ circuit evaluation problem, which is complete for this class under $\mathsf{L}$ reductions. This establishes the main result.
Overall, our results provide an exact characterization of the expressive power of padded fixed-depth transformers as $\mathsf{TC}^{0}$, answering an open question (Pfau et al., 2024). Further, we exactly characterize the expressive power of polylog-looped padded transformers, which reveals that padding and looping dramatically extend the expressive power of transformers under standard complexity conjectures. We take these results to motivate empirical investigation of padding and looping as forms of inference-time compute that are more parallelizable than standard CoT.
[…]
Figure 1: Summary of our core results: exact charactizations of the expressive power of padded fixed-depth transformers and $O(\log^{d}n)$-depth looped, padded transformers. Via Theorem 1, we show that $\mathsf{AHAT}^{0}_{*}=\mathsf{FO}\textrm{-uniform}\;\mathsf{TC}^{0}$. Combining Theorem 1 with Theorems 2 and 3, we show that, for $d\geq 1$, $\mathsf{AHAT}^{d}_{*}=\mathsf{L}\textrm{-uniform}\;\mathsf{TC}^{d}$. More finegrained results are discussed in each section.
2 Preliminaries
2.1 Averaging-Hard-Attention Transformers
Following previous work, we analyze a model of transformers that is slightly idealized compared to standard soft-attention transformers. Specifically, we analyze transformers with averaging hard attention (AHAT; also called “saturated” by Merrill et al., 2022) and masked pre-norm (Merrill and Sabharwal, 2023b). AHAT attention heads only attend to the values that maximize the attention score; in the case of ties, the head returns a uniform average of all values with tied scores. Masked pre-norm means that the transformer can apply a linear projection before pre-norm at the beginning of each sublayer, which is useful for reading individual values from the residual stream without interference from other values. Overall, we believe these idealizations are minimal changes that make it easier to implement algorithmic constructions that generalize to any sequence length.
More formally, we define the transformer sublayers in this AHAT model following Merrill and Sabharwal (2024). Each sublayer uses masked pre-norm (Xiong et al., 2020; Merrill and Sabharwal, 2023b), reading input $\mathbf{z}_{i}=\mathsf{layer\_norm}(\mathbf{m}\mathbf{h}_{i})$, where $\mathbf{h}_{i}$ is the previous sublayer output and layer-norm can be standard layer-norm (Ba et al., 2016) or RMS norm (Zhang and Sennrich, 2019). The sublayer then outputs $\delta_{1},\ldots,\delta_{n}$, and the residual stream is updated as $\mathbf{h}^{\prime}_{i}=\mathbf{h}_{i}+\delta_{i}$.
Definition 1 (Self-attention sublayer).
The self-attention sublayer is parameterized by a mask $\mathbf{m}\in\mathbb{Q}^{m}$, output projection matrix $\mathbf{W}:\mathbb{Q}^{m}\to\mathbb{Q}^{m}$, and, for $1\leq k\leq h$, query, key, and value matrices $\mathbf{Q}^{k},\mathbf{K}^{k},\mathbf{V}^{k}$, each of which is a projection from $\mathbb{Q}^{m}$ to $\mathbb{Q}^{m/h}$.
Given input $\mathbf{z}_{i}$, the self-attention sublayer computes queries $\mathbf{q}_{i}=\mathbf{Q}^{k}\mathbf{z}_{i}$, keys $\mathbf{k}_{i}=\mathbf{K}^{k}\mathbf{z}_{i}$, and values $\mathbf{v}_{i}=\mathbf{V}^{k}\mathbf{z}_{i}$. Next, these values are used to compute the attention head outputs:
$\mathbf{a}_{i,k}=\lim_{\tau\to 0}\sum_{j=1}^{c}\frac{\exp(1/\tau\cdot\mathbf{q}_{i,k}^{\top}\mathbf{k}_{j,k})}{Z_{i,k}}\cdot\mathbf{v}_{j,k},\quad\textrm{where}\;Z_{i,k}=\sum_{j=1}^{c}\exp\left(1/\tau\cdot\mathbf{q}_{i,k}^{\top}\mathbf{k}_{j,k}\right).$
We can set $c=i$ to define causally masked attention and $c=n$ for unmasked attention. Averaging hard attention is formalized by taking the low-temperature limit ($\tau\to 0$), which causes all probability mass to be concentrated on the tokens that maximize the attention score. In practice, transformers can approximate this by learning a temperature close to zero; for a fixed sequence length, this approximation will hold, but for longer strings it will break down. Finally, the output of the self-attention sublayer is computed by aggregating the head outputs via $\delta_{i}=\mathbf{W}\cdot\mathrm{concat}(\mathbf{a}_{i,1},\ldots,\mathbf{a}_{i,h})$.
Definition 2 (Feedforward sublayer).
The feedforward sublayer at layer $\ell$ is parameterized by a mask $\mathbf{m}\in\mathbb{Q}^{m}$ and projections $\mathbf{W}:\mathbb{Q}^{m}\to\mathbb{Q}^{w}$ and $\mathbf{U}:\mathbb{Q}^{w}\to\mathbb{Q}^{m}$.
A feedforward layer computes a local update to the residual stream via $\delta_{i}=\mathbf{U}\cdot\mathsf{ReLU}(\mathbf{W}\mathbf{z}_{i})$.
A transformer defines a function $\Sigma^{*}\to\Sigma$ or $\Sigma^{*}\to\{0,1\}$ if we add a linear head to the final layer and take its argmax as the output. We say that a transformer $T:\Sigma\to\{0,1\}$ recognizes a language $L$ (with beginning-of-sequence token \$currency-dollar\\$\$) if, for any $w\in\Sigma^{*}$, $T(\$w)=1$ if and only if $w\in L$.
Precision.
In practice, transformers use low precision (relative to $\lvert w\rvert$) during training and inference. To justify that our constructions do not unrealistically rely on too much precision, we make sure that they use at most $O(c\log n)$ precision, following previous work (Merrill and Sabharwal, 2023a).
Layer-Norm Hash.
Our constructions will use the layer-norm hash representation (Merrill and Sabharwal, 2023b, 2024) for query-key matching with positive integer values.
Definition 3.
Given $z\in\mathbb{N}$, its layer-norm hash is the vector $\langle z,1,-z,-1\rangle/\sqrt{z^{2}+1}$
The layer-norm hash is computable within a transformer and satisfies that property that $\phi(i)^{\top}\cdot\phi(j)=1$ if and only if $i=j$. This makes it useful for retrieval based on position matching using AHATs.
2.2 Padded and Looped AHATs
We assume a looped transformer model as previously defined by Merrill and Sabharwal (2024):
Definition 4 ($d(n)$-looped transformer).
A looped transformer’s layers are partitioned into blocks $\langle A,B,C\rangle$. On an input of length $n$, block $B$ is repeated depth-wise $d(n)$ times.
Looped transformers provide a way of dynamically scaling width at test time by looping a block of layers. In addition, we also consider padding (Pfau et al., 2024) as a simple way to increase width:
Definition 5 ($w(n)$-padded transformer).
On an input of length $n$, we first append $w(n)$ “blank” tokens ($\square\not\in\Sigma$) and then feed this full string through the transformer.
Looping and padding can also be combined, giving us the following language class for transformers:
Definition 6 (Padded and looped transformers).
Let $d,k\in\mathbb{Z}_{\geq 0}$. $\mathsf{AHAT}^{d}_{k}$ is the class of languages recognizable by a causally masked universal transformer with masked pre-norm, a beginning-of-sequence token, no position embedding, $O(\log^{d}n)$ depth, and $O(n^{k})$ padding tokens. Further, $\mathsf{AHAT}^{d}_{*}=\bigcup_{k=0}^{\infty}\mathsf{AHAT}^{d}_{k}$, $\mathsf{AHAT}^{*}_{k}=\bigcup_{d=0}^{\infty}\mathsf{AHAT}^{d}_{k}$, and $\mathsf{AHAT}^{*}_{*}=\bigcup_{d=0}^{\infty}\mathsf{AHAT}^{d}_{*}$.
$\mathsf{AHAT}^{d}_{k}$ uses causal masking with no positional embeddings. In contrast, some of our constructions will be for unmasked transformers with $i/n$ position embeddings (standard in theoretical analysis; Merrill and Sabharwal, 2023a). We will write $\mathsf{uAHAT}^{d}_{k}$ to signify these classes. Similarly, we will occasionally consider mixed-masked transformers where some attention heads use causal masking and others do not. We will denote these classes as $\mathsf{mAHAT}^{d}_{k}$.
With some abuse of notation, we will use $\mathsf{AHAT}^{d}_{k}$ to denote both the class of languages as defined above as well as the corresponding class of transformer models, i.e., transformers with $O(\log^{d}n)$ and $(n^{k})$ padding tokens. The distinction should be clear from the context.
2.3 Circuit Complexity
We define the circuit complexity classes $\mathsf{AC}^{d}$ and $\mathsf{TC}^{d}$ in the standard way (Arora and Barak, 2009; Strobl et al., 2024). A circuit family is a mapping $\mathcal{C}=\{C_{n}\}_{n=0}^{\infty}$ from input lengths $n$ to circuits that take $n$ inputs. $\mathsf{AC}^{d}$ is the class of language that can be recognized by polynomial-size, fixed-depth circuit families with unbounded-arity AND/OR gates. $\mathsf{TC}^{d}$ is the same class but augmented with unbounded-arity MAJ gates that take a majority vote over their input bits.
It will often be useful to talk about uniform variants of these classes. A $\mathsf{X}$-uniform circuit family obeys that constraint that circuit $C_{n}$ for input size $n$ can be “built” from the input string $1^{n}$ using computation in $\mathsf{X}$ (For a more formal definition, see Strobl et al., 2024). We will consider two standard types of uniformity conditions: $\mathsf{FO}$ uniformity (which is equivalent to $\mathsf{DLOGTIME}$ uniformity; Mix Barrington et al., 1990) and the weaker $\mathsf{L}$ uniformity (i.e., log-space uniformity).
Finally, we will also use the notion of completeness for circuit classes. Informally, an $\mathsf{X}$-complete problem is a problem in $\mathsf{X}$ to which any other problem in $\mathsf{X}$ can be mapped via some simple reduction $\mathsf{R}$. Whether a problem is complete depends on the notion of reduction used. See Definition 9 for a more formal definition of reductions, which makes completeness fully defined.
2.4 Logic
We define the standard notion first-order logic over strings ($\mathsf{FO}$; cf. Merrill and Sabharwal, 2023c). $\mathsf{FO}$ formulas map strings to boolean values. Its formulas can check for token occurrences at specific string positions, and it allows quantification over positions in the input string.
Definition 7.
More formally, $\mathsf{FO}$ can be defined as follows. $\mathsf{FO}$ contains two types: indices, representing positions in the input string, and formulas, which evaluate to true or false. For an index or formula $x$, we write $\llbracket x\rrbracket^{w,v}$ for the evaluation of $x$ on string $w$ with variable assignments $v$ (a map from names to values). Indices in $\mathsf{FO}$ are integers denoting positions in the input string:
  1. 1.
    The constant $1$, representing the first token’s position: $\llbracket 1\rrbracket^{w,v}=1$.
  2. 2.
    The constant $n$, representing the last token’s position: $\llbracket n\rrbracket^{w,v}=\lvert w\rvert$.
  3. 3.
    Symbols (e.g., $i,j,k$) representing variables ranging over positions $1$ to $n$: $\llbracket i\rrbracket^{w,v}=v[i]$.
Formulas in $\mathsf{FO}$ are then constructed as follows:
  1. 1.
    Let $\Sigma$ be a finite alphabet. For each $\sigma\in\Sigma$ and any index $i$, $Q_{\sigma}(i)$ is a formula that is true when the $i$-th input token is $\sigma$. That is, $\llbracket Q_{\sigma}(i)\rrbracket^{w,v}=1$ iff $w_{m}=\sigma$ where $m=\llbracket i\rrbracket^{w,v}$.
  2. 2.
    For two indices $i,j$, $i=j$, $i\leq j$, and $i\geq j$ are formulas with their conventional semantics.
  3. 3.
    For two indices $i,j$, $\mathsf{bit}(i,j)$ is a formula returning the $j$-th bit of $i$.
  4. 4.
    For two formulas $P,Q$, $P\wedge Q$ and $P\vee Q$ are formulas with their conventional semantics.
  5. 5.
    For any formula $P$ (which may refer to $i$ or any over variable), the following are formulas:
    1. (a)
      $\exists i.P$ means setting $i$ to some value $m\in[1,n]$ makes $P$ true (more formally, $\llbracket P\rrbracket^{w,v\mid i=m}=1$).
    2. (b)
      $\forall i.P$ means setting $i$ to any value $m\in[1,n]$ make $\phi$ true.
An $\mathsf{FO}$ formula $P$ with no free variables is called a sentence and returns a value in $\{0,1\}$ for each input string. The language defined by a sentence is the set of strings mapped to 1. The nesting depth of an $\mathsf{FO}$ formula is the depth of its syntactic tree constructed by the rules above. The number of distinct variables is the number of different variable names used in $P$. This can be minimized by allowing two independent quantifiers in parallel subformulas to use the same variable name.
It is known that the languages defined by $\mathsf{FO}$ are exactly the circuit complexity class $\mathsf{DLOGTIME}$-uniform $\mathsf{AC}^{0}$ (Mix Barrington et al., 1990). This class captures hard-attention transformers (Hao et al., 2022; Yang et al., 2024), but it is not large enough to capture soft-attention transformers (Merrill and Sabharwal, 2023a). To consider a more expressive logic capable of modeling soft-attention transformers, we can extend $\mathsf{FO}$ with majority quantifiers (Merrill and Sabharwal, 2023c). Specifically, we will define $\mathsf{FO}[\mathsf{M}^{2}]$ as $\mathsf{FO}$ extended to include a paired majority quantifier:
Definition 8 (Mix Barrington et al., 1990).
$\mathsf{FO}[\mathsf{M}^{2}]$ is $\mathsf{FO}$ extended to include the $\mathsf{M}^{2}$ quantifier, defined as follows: $\mathsf{M}^{2}(i,j).P(i,j)$ is true if $\phi(i,j)$ holds for a majority of pairs of positions $(i,j)\in[n]^{2}$ in the string.
It is known that $\mathsf{FO}[\mathsf{M}^{2}]$ defines exactly $\mathsf{DLOGTIME}$-uniform $\mathsf{TC}^{0}$ (Mix Barrington et al., 1990). It is also possible to use majority quantifiers to define addition over indices, so without loss of generality, we can assume $\mathsf{FO}[\mathsf{M}^{2}]$ formulas have no addition (unlike $\mathsf{FO}$ formulas). Moreover, it is possible to simulate $\mathsf{bit}$ in terms of $\mathsf{M}^{2}$, so we can consider $\mathsf{FO}[\mathsf{M}^{2}]$ formulas not to contain $\mathsf{bit}$, in contrast to the more standard logic $\mathsf{FO}[\mathsf{M},\mathsf{bit}]$ that also defines $\mathsf{TC}^{0}$. This makes $\mathsf{FO}[\mathsf{M}^{2}]$ a simpler target for our transformer constructions than $\mathsf{FO}[\mathsf{M},\mathsf{bit}]$.
3 Masked and Unmasked Transformers
Before presenting our main results, we briefly discuss a simple but important relationship between different kinds of masking that will come in handy. When using a transformer as a recognizer for formal languages, there is a choice of whether to use an encoder (with no attention masking) or a decoder (with causal attention masking) to encode the input string. Causally masked (decoder) models are more standard these days, as well as theoretically more convenient (e.g., no need for position embedding, as position information can be derived from causal attention), so we take them as our default model. However, it is sometimes easier to reason about how to solve problems with unmasked transformers. Fortunately, the following lemma shows we can simulate an unmasked transformer (encoder) with a causally masked transformer (decoder), if we allow padding tokens proportional to the transformer depth. This will be useful going forward in several places where we convert unmasked and mixed-masked constructions to causally masked constructions.
Lemma 1 (Unmasked to Causally Masked).
Let $E$ be an unmasked AHAT encoder with depth $\ell\geq 1$. Then there exists a causally masked AHAT decoder $D$ with depth $\ell$ and with $\ell n$ padding tokens on input length $n$ that is equivalent to $E$ in the following sense: the final $n$ outputs of $D$ match the original $n$ outputs of $E$.
Proof.
The idea is that $D$ will unroll a sequence of $\ell$ “blocks” on the padding length dimension, each of width $n$. Each block will attend to the previous block in the previous layer and can thus see all tokens despite causal masking.
To implement the block construction, we first compute $\phi(n)$. We then use this to compute $\phi(b_{i})$, where $b_{i}=\lfloor i/n\rfloor\in[1,\ell]$, which represents the block that each token belongs to. We also compute $\phi(b_{i}-1)\in[0,\ell-1]$ using a head that ignores the bos token. Next, we modify each original head from $U$ to have an additional term in the query/key product. The query is $C\phi(b_{i}-1)$ and the key is $\phi(b_{j})$, where $C$ is a large weight. We set $C$ to a fixed large value (independent of $n$) such that this term dominates all other terms in the inner product computation at each token. As a result, the head is constrained over keys in the context where $b_{j}=b_{i}-1$, i.e., the keys from the last block. Within these keys, it computes exactly the same AHAT output as the original masked unmasked head. In this way, we are able to simulate a causally masked transformer with an unmasked transformer. ∎
This allows us to relate various masking variants of $\mathsf{AHAT}$ classes as follows:
Proposition 1.
Unmasked and mixed-masked padded transformers (i.e., uAHATs and mAHATs) can be simulated by causally masked transformers (AHATs). Specifically, for any $d,k\in\mathbb{N}$, the following holds for the corresponding problem classes:
  1. 1.
    $\mathsf{uAHAT}^{0}_{k}\subseteq\mathsf{mAHAT}^{0}_{k}\subseteq\mathsf{AHAT}^{0}_{\max\{k,1\}}$
  2. 2.
    $\mathsf{uAHAT}^{d}_{k}\subseteq\mathsf{mAHAT}^{d}_{k}\subseteq\mathsf{AHAT}^{d}_{1+\max\{k,1\}}$ for $d\geq 1$
  3. 3.
    $\mathsf{uAHAT}^{d}_{*}\subseteq\mathsf{mAHAT}^{d}_{*}\subseteq\mathsf{AHAT}^{d}_{*}$.
We defer the proof of this to Appendix A.
4 Fixed-Depth Padded Transformers Recognize Exactly $\mathsf{TC}^{0}$
We are now ready to prove our first main result, namely that padding tokens allow transformers to simulate $\mathsf{FO}$ formulas—with more padding allowing more nesting of variables. Moreover, they can simulate formulas with the special paired majority quantifiers.
Lemma 2.
An $\mathsf{FO}[\mathsf{M}^{2}]$ formula with $k\geq 1$ distinct variables and nesting depth $\ell$ can be computed in $\mathsf{uAHAT}^{0}_{k}$ (and hence in $\mathsf{AHAT}^{0}_{k}$) with (fixed) depth $\ell$.
Proof.
We will store all possible configurations of $k$ variables using $n^{k}$ padding tokens. Each token corresponds to a specific assignment of all the variables, which we denote $v$ (cf. Definition 8). We will present an inductive transformer construction that uses a single layer to compute the boolean variable of each formula and the integer value of each numerical expression in $\mathsf{FO}[\mathsf{M}^{2}]$, assuming the constituent formulas and values were already computed at the previous layer. Since each $\mathsf{FO}[\mathsf{M}^{2}]$ formula consists of a finite number of subformulas and values, we can accumulate all constituents in the residual stream in order to compute any larger formula.
In more detail, let $\llbracket x\rrbracket^{w,v}$ be the value of a formula or numerical value $x$ on string $w$ under assignment $v$. We identify each padding token $v$ with a variable assignment $v$ so that, for every $P$, token $v$ stores $\llbracket P\rrbracket^{w,v}$ in the residual stream with a scalar whose sign indicates a truth value. For a numerical value $i$, token $v$ will represent $\llbracket i\rrbracket^{w,v}$ as a small vector $\phi(\llbracket i\rrbracket^{w,v})$. where $\phi$ is the layer-norm hash (Definition 3). We show how to evaluate each constituent of an $\mathsf{FO}[\mathsf{M}^{2}]$ formula. As mentioned after Definition 8, addition over indices and $\mathsf{bit}$ are subsumed by $\mathsf{M}^{2}$, so we do not need to simulate them.
  1. 1.
    Variables. At each token $v$, we will compute $\phi(\llbracket i\rrbracket^{w,v})$, a representation of the value of variable $i$ under assignment $v$. We view the integer $v$ as a tuple of $k$ values in $[1,n]$, so we will take a “projection” to retrieve $v[i]$ using quotient and remainder operations (Merrill and Sabharwal, 2024, Lemma 3.1). More formally, if $i$ is the $m$-th variable, compute
    $\phi(\lfloor v/n^{m-1}\rfloor\bmod n)=\phi(\llbracket i\rrbracket^{w,v}).$
    Thus, we conclude that we can compute $\phi(\llbracket i\rrbracket^{w,v})$ at each token $v$.
  2. 2.
    Constants. To compute the constant $1$ or $n$ at each configuration $v$, we can attend from each $v$ to the first and last input token and retrieve its position to get $\phi(1)$ or $\phi(\lvert w\rvert)$.
  3. 3.
    Comparisons. Given two numerical values $i,j$ (either variables or constants) already stored at $v$ as $\phi(\llbracket i\rrbracket^{w,v}),\phi(\llbracket j\rrbracket^{w,v})$, we can simply compare $\phi(\llbracket i\rrbracket^{w,v})-\phi(\llbracket j\rrbracket^{w,V})$ on some axis and apply layer-norm to obtain $\llbracket i>j\rrbracket^{w,v}$ at token $v$.
  4. 4.
    Token Predicates. Assume we have $\phi(\llbracket i\rrbracket^{w,v})$ previously stored at assignment $v$. We can hard-attend to retrieve token $w_{m}$ where $m=\llbracket i\rrbracket^{w,v}$. We then compute $\llbracket Q_{\sigma}(i)\rrbracket^{w,v}$ by checking whether $w_{m}=\sigma$.
  5. 5.
    Connectives. Assume we have $\llbracket P\rrbracket^{w,v}$ and $\llbracket Q\rrbracket^{w,v}$ previously stored at each assignment token $v$. Then we can simply use a feedforward network to compute $\llbracket\neg P\rrbracket^{w,v}$, $\llbracket P\wedge Q\rrbracket^{w,v}$, or $\llbracket P\vee Q\rrbracket^{w,v}$ independently at each $v$.
  6. 6.
    Standard Quantifiers. Assume we have $\llbracket P\rrbracket^{w,v}$ stored at each configuration $v$. Then, at each $m$, we want to resolve the quantifier $\mathsf{Q}$ over the set $\{\llbracket P\rrbracket^{w,v|i=m}\}_{m=1}^{n}$, where $v|i=m$ denotes $v$ with $i$ overriden to have value $m$. We will count $c$, the number of $m$ such that $P^{v|i=m}$ holds. More formally, let $j_{1},\ldots,j_{k-1}$ be the set of variables excluding $i$. Then we can attend from $v$ over $v^{\prime}$ with query $\langle\phi(j_{1}^{v},\ldots,\phi(j_{k-1}^{v}\rangle$, key $\langle\phi(j_{1}^{v^{\prime}},\ldots,\phi(j_{k-1}^{v^{\prime}})\rangle$, and value $P^{v^{\prime}}$ to retrieve $c/n$. Finally, we threshold $c/n$ against $\frac{1}{2n}$ (for $\exists$) or $\frac{2n-1}{2n}$ (for $\forall$) to resolve $(\mathsf{Q}i.\phi)^{v}$.
  7. 7.
    Paired Majority Quantifiers. Assume we have $P^{v}$ stored at each configuration $v$. We want to compute $(\mathsf{M}(i,j).P)^{v}$ for any two variables $i,j$ already represented. The idea slightly generalizes the construction for single-variable quantifiers: we will use attention to count $c$, the number of assignments where $P^{v|i=m,j=\ell}$ is true. Formally, let $j_{1},\ldots,j_{k-2}$ be the set of variables excluding $i,j$. Then we can attend from $v$ over $v^{\prime}$ with query $\langle\phi(j_{1}^{v},\ldots,\phi(j_{k-2}^{v}\rangle$, key $\langle\phi(j_{1}^{v^{\prime}},\ldots,\phi(j_{k-2}^{v^{\prime}})\rangle$, and value $P^{v^{\prime}}$ to retrieve $c/n^{2}$. Finally, we threshold $c/n^{2}$ against $\frac{1}{2}$ to resolve $(\mathsf{M}^{2}(i,j).P)^{v}$.
In conclusion, we can compute any $\mathsf{FO}[\mathsf{M}^{2}]$ formula in $\mathsf{uAHAT}^{0}_{k}$ by inductively computing its constituent formulas and storing their values over $k$ variables using $n^{k}$ padding tokens. The result extends to $\mathsf{AHAT}^{0}_{k}$ by applying Proposition 1. ∎
Our construction for Lemma 2 somewhat resembles (though is distinct from) a result of Lange (2004, Corollary 6.8) that any problem in $\mathsf{TC}^{0}$ can be reduced to majority logic via a transformation that appends some extra tokens to the input. However, their transformation is not padding since it appends some “non-blank” tokens, and their result also does not clearly apply to transformers.
Combined with previous results about transformers being in $\mathsf{FO}$-uniform $\mathsf{TC}^{0}$, Lemma 2 yields an exact characterization of constant-depth transformers with padding:
Theorem 1.
$\mathsf{uAHAT}^{0}_{*}=\mathsf{AHAT}^{0}_{*}=\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}$, i.e., $\mathsf{FO}[\mathsf{M},\mathsf{bit}]$.
Proof.
It is known that $\mathsf{AHAT}^{0}\subseteq\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}$ (Merrill and Sabharwal, 2023a, c; Chiang, 2025), and this generalizes to transformers with padding tokens (Pfau et al., 2024). We will show $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}\subseteq\mathsf{uAHAT}^{0}_{*}$. Mix Barrington et al. (1990, Proposition 10.3) proved that $\mathsf{FO}$-uniform $\mathsf{TC}^{0}$ is definable by $\mathsf{FO}$ formulas with $\mathsf{M}^{2}$ quantifiers: notably, the $\mathsf{bit}$ predicate is not necessary when using $\mathsf{M}^{2}$ quantifiers. Thus, Lemma 2 establishes that $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}\subseteq\mathsf{uAHAT}^{0}_{*}$. Finally, from Proposition 1, we have $\mathsf{uAHAT}^{0}_{*}\subseteq\mathsf{AHAT}^{0}_{*}$. Hence, $\mathsf{uAHAT}^{0}_{*}=\mathsf{AHAT}^{0}_{*}=\mathsf{FO}$-uniform $\mathsf{TC}^{0}$. ∎
A technical hurdle to obtaining this characterization in prior work was simulating the $\mathsf{bit}$ predicate in standard definitions of $\mathsf{TC}^{0}$ (Pfau et al., 2024). Our results circumvent this by instead relating transformers to $\mathsf{FO}[\mathsf{M}^{2}]$, which is equivalent to standard $\mathsf{FO}[\mathsf{M},\mathsf{bit}]$ (Mix Barrington et al., 1990). It thus follows from Theorem 1 that padded transformers can also simulate $\mathsf{bit}$ as well as all of $\mathsf{FO}$, which will be useful in the following section for simulating $\mathsf{FO}$ reductions.
5 Logd-Looped Padded Transformers Recognize Exactly $\mathsf{TC}^{d}$
The notion of completeness of a problem (or language) for a complexity class under certain types of reduction has played a key role in computational complexity. Just like it has been useful for reasoning about the expressivity of resource-bounded standard models of computation (Turing machines, circuit models, etc.), we show it can also be used to reason about the expressivity of padded transformers.
We begin by formally defining the notion of reductions in terms of predicates or languages (rather than string-to-string functions). This will make it easier to precisely implement reductions inside transformers, which produce contextual representations of prefixes of the input string in parallel, in contrast to the more standard definition of reductions as string-to-string functions.
Definition 9.
Let $b(i)$ be the binary encoding of $i\in\mathbb{N}$ in some alphabet $\Sigma\supseteq\{0,1\}$. Let $\mathsf{R}$ be a class of languages. We say a transduction $f:\Sigma^{*}\to\Sigma^{*}$ is an $\mathsf{R}$ reduction if $\lvert f(w)\rvert$ is polynomial in $\lvert w\rvert$ and the language $R_{f}=\{(w,b(i),\sigma)\mid f(w)_{i}=\sigma\}$ is in $\mathsf{R}$ [^1] Here $f(w)_{i}$ denotes the $i$-th token of $f(w)$ if $i\leq\lvert f(w)\rvert$, and a special symbol $\square\not\in\Sigma$ otherwise.
Definition 9 recovers the standard notions of $\mathsf{FO}$ and $\mathsf{L}$ reductions. A transformer can be said to compute a reduction $f$ if it recognizes the language of triples $(w,i,\sigma)$ defined above. Since $\lvert\Sigma\rvert$ is finite and transformers can equally easily output a ‘token’ in $\Sigma$ instead of just 0/1, it is in fact more natural to require the transformer to compute a functional form of this language, namely compute $r_{f}(w,i)$ defined as $f(w)_{i}$. Our constructions work under both views, though the latter is often more natural and efficient. Formally:
Definition 10.
We say that a transformer computes an $\mathsf{R}$ reduction $f$ if it either recognizes the language $R_{f}=\{(w,i,\sigma)\mid f(w)_{i}=\sigma\}$ in $\mathsf{R}$ or computes the function $r:\Sigma\times\mathbb{N}\to\Sigma$ defined as $r_{f}(w,i)=f(w)_{i}$, where, in either case, $i$ is encoded in binary.
Lemma 3.
Let $\mathsf{C},\mathsf{R}$ be classes of languages. Let language $L$ be $\mathsf{C}$-complete under $\mathsf{R}$ reductions. If $\mathsf{AHAT}^{d}_{*}$ transformers can recognize $L$ and compute every $\mathsf{R}$ reduction, then $\mathsf{C}\subseteq\mathsf{AHAT}^{d}_{*}$.
Proof.
Consider any language $L^{\prime}\in\mathsf{C}$. By the assumed completeness of $L$, there exists an $\mathsf{R}$ reduction $f$ that maps inputs of $L^{\prime}$ into inputs of $L$ such that $w\in L^{\prime}$ if and only if $f(w)\in L$. From the last precondition of the theorem, there exists a causally masked logd-depth padded transformer $T_{f}$ that recognizes the corresponding reduction language $R_{f}$ or, equivalently, computes the reduction function $r_{f}$ (cf. Definition 10). We will assume the latter, i.e., that $T_{f}$ computes $r_{f}$, though the construction can also be made to work if $T_{f}$ checks membership in $R_{f}$. Additionally, we also have from a precondition that there exists a causally masked logd-depth padded transformer $T_{L}$ that recognizes $L$.
The idea is to “stack” $T_{L}$ on top of $T_{f}$ to obtain a logd-depth padded transformer $T$ that recognizes $L^{\prime}$. Intuitively, given an input $w$, the first set of layers of $T$ will compute $f(w)$ tokenwise, by computing $r(w,i)=f(w)_{i}$ for every $i$ in parallel. To this end, we will essentially make $\lvert f(w)\rvert$ copies of the padding tokens needed by $T_{f}$ and perform the computation of $T_{f}$ independently for each $f(w)_{i}$. The second and final set of layers of $T$ will then check whether the string $f(w)$ produced by the first set of layers is in $L$, which will hold if and only if $w\in L^{\prime}$. We next make this idea more concrete.
Compute Reduction. Suppose $T_{f}\in\mathsf{AHAT}^{d}_{k}$ and let $n=\lvert w\rvert$. Then, on input $(w,i)$, where $i$ is represented in binary, $T_{f}$ uses $O(n^{k})$ padding tokens to compute $f(w)_{i}$, which we upper bound by $n^{k+1}$ for sufficiently large $n$ (for small $n$, we assume $T_{f}$ instead uses a finite lookup table (and no padding). There is a uniform $\mathsf{TC}^{0}$ circuit family that computes $\mathsf{bit}(i,j)$ from input $\phi(i),\phi(j)$, so, by Theorem 1, there exists a transformer $T_{\mathsf{bit}}$ that computes $\mathsf{bit}$ with $n^{b}$ padding tokens, for some $b$ and sufficiently large $n$ (again, for smaller $n$, we assume $T_{\mathsf{bit}}$ uses a finite lookup table). Since $\lvert f(w)\rvert$ is bounded by some polynomial $n^{c}$, we know that any $i\leq\lvert f(w)\rvert$ takes at most $c\log n$ bits to specify, which we upper bound by $c^{\prime}=n$ (since it’s unclear how a transformer would compute $\log n$).
We will construct $n^{c}$ blocks of padding tokens, each of size $B=c^{\prime}n^{b}+n^{k+1}$. Block $i$ will first compute the binary expansion of $i$ in its first $c^{\prime}n^{b}$ padding tokens using $T_{\mathsf{bit}}$, and then use the remaining $n^{k+1}$ padding tokens to consume $(w,i)$ using $T_{f}$ and return $r_{f}(w,i)$. At each token $t$ in block $i$, we compute $\phi(i)$ as $\phi(\lfloor t/B\rfloor)$, using the fact that we can compute integer division with a fixed block of transformer layers (Merrill and Sabharwal, 2024) and having computed $B$ as a function of $n$ in earlier layers (since it is in $\mathsf{TC}^{0}$). For $t^{\prime}\leq c^{\prime}$, token $t=t^{\prime}n^{b}$ computes $\phi(t^{\prime})$ via division similarly to $\phi(i)$ and then $\mathsf{bit}(i,t^{\prime})$. This recovers the binary representation of $i$ stored across tokens $t=t^{\prime}n^{b}$ in block $i$, for $1\leq t^{\prime}\leq c^{\prime}$. Finally, we apply a slightly modified $T_{f}$ over the block. Specifically, we add a new term to each attention head so that it only attends over the input tokens and some tokens within the current block: those satisfying $t=t^{\prime}n^{b}$ for some $1\leq t^{\prime}\leq c^{\prime}$ or $t>c^{\prime}n^{b}$: since $c^{\prime}=n$, these predicates are simple to check. As a result, block $i$ simulates $T_{f}$ over $w$ with $n^{k}$ padding. Thus, the final token of block $i$ computes $r_{f}(w,i)=f(w)_{i}$.
Solve Complete Problem. We will use additional layers to check whether $f(w)\in L^{\prime}$. We are given that there exists a transformer $T_{L}$ that, on input $w^{\prime}$, checks whether $w^{\prime}\in L^{\prime}$. For each attention head in $T_{L}$, we add a new term to the attention score that is very negative if that token is not the final token in some attention block. Thus, each head in $T_{L}$ will only attend over tokens that are final in some block. We also modify $T_{L}$ so that it uses $\phi(i)$ in place of the position embedding for token $t$. Thus, $T_{L}$ computes whether $f(w)\in L^{\prime}$, which is equivalent to recognizing whether $w\in L$. ∎
Combined with results in prior work (see proof below), it follows from Lemma 3 that padded log-depth transformers can recognize any language in $\mathsf{NL}$:
Theorem 2.
$\mathsf{NL}\subseteq\mathsf{AHAT}^{1}_{*}$.
Proof.
Let $L$ be the graph connectivity problem, class $\mathsf{C}$ be $\mathsf{NL}$, and class $\mathsf{R}$ be $\mathsf{FO}$. We will show that the preconditions of Lemma 3 are met, from which it will follow that $\mathsf{NL}\subseteq\mathsf{AHAT}^{1}_{*}$. First, graph connectivity is known to be $\mathsf{NL}$-complete under $\mathsf{FO}$ reductions (Immerman, 1998). Second, Merrill and Sabharwal (2024) recently showed that mixed-masked log-depth transformers with cubic padding can recognize the graph connectivity problem $L$, i.e., $L\in\mathsf{mAHAT}^{1}_{*}$. From Proposition 1, it follows that $L\in\mathsf{AHAT}^{1}_{*}$. Finally, it follows from Theorem 1 that fixed-depth causally masked transformers can recognize languages in $\mathsf{FO}[\mathsf{M},\mathsf{bit}]$, and hence also languages in $\mathsf{FO}$. Such transformers can therefore compute $\mathsf{FO}$ reductions $f$ in the sense of Definition 10, i.e., compute the function $r_{f}(w,i)$ defined as $f(w)_{i}$, the $i$-th bit of $f(w)$. Thus, Lemma 3 applies. ∎
Since $\mathsf{L}$ reductions are in $\mathsf{NL}$, we can bootstrap this result to obtain the following stronger result. For this, we will leverage completeness of the problem of evaluating a given logd depth circuit, formalized in Definition 11 (Appendix B) as the $\mathsf{TC}^{d}$ circuit evaluation problem. We first note the following extension of a known result about fixed-depth transformers (see Appendix A for a proof):
Lemma 4.
For any $d\geq 1$, $\mathsf{AHAT}^{d}_{*}\subseteq\mathsf{L}\textrm{-uniform}\ \mathsf{TC}^{d}$.
In contrast to Theorem 1, Lemma 4 applies with $\mathsf{L}$ uniformity instead of $\mathsf{FO}$ uniformity. We will show the other direction of Lemma 4 to establish $\mathsf{L}$-uniform $\mathsf{TC}^{d}$ as an exact characterization of $\mathsf{AHAT}^{d}_{*}$:
Theorem 3.
For any $d\geq 1$, $\mathsf{AHAT}^{d}_{*}=\mathsf{L}$-uniform $\mathsf{TC}^{d}$. [^2] Theorem 3 generalizes to $\mathsf{NL}$-uniform $\mathsf{TC}^{d}=\mathsf{AHAT}^{d}_{*}$, which implies $\mathsf{NL}$-uniform $\mathsf{TC}^{d}=\mathsf{L}$-uniform $\mathsf{TC}^{d}$. At first glance, it might appear suspicious that we can prove this complexity result through our analysis of transformers. However, we verify in Appendix C via a simpler proof that this equivalence indeed holds.
Proof.
Lemma 4 directly provides one direction. To show the other direction that $\mathsf{L}$-uniform $\mathsf{TC}^{d}\subseteq\mathsf{AHAT}^{d}_{*}$, we will apply Lemma 3 with the wide-$\mathsf{TC}^{d}$ circuit evaluation problem (Appendix B) as $L$, $\mathsf{L}$-uniform $\mathsf{TC}^{d}$ as class $\mathsf{C}$, and $\mathsf{L}$ as class $\mathsf{R}$. We next argue that the preconditions of Lemma 3 are met, finishing the proof. First, 7.1 (Appendix B) shows that logd-depth looped transformers (without padding) can solve the wide-$\mathsf{TC}^{d}$ circuit evaluation problem, i.e., $L\in\mathsf{AHAT}^{d}_{0}$. Combining this with Lemma 4, we also obtain that $L\in\mathsf{L}\textrm{-uniform}\ \mathsf{TC}^{d}$. Lemma 6 (Appendix B) shows that $L$ is hard for $\mathsf{L}$-uniform $\mathsf{TC}^{d}$ under $\mathsf{L}$ reductions. Together, we obtain that $L$ is complete for this class. Finally, Theorem 2 implies that logd-depth transformers can recognize any language in $\mathsf{L}$, and thus compute any $\mathsf{L}$ reduction in the sense of Definition 10. Applying Lemma 3, we conclude that $\mathsf{L}$-uniform $\mathsf{TC}^{d}\subseteq\mathsf{AHAT}^{d}_{*}$. ∎
6 Conclusion
Our results in this work give a precise theoretical understanding of how padding and looping—two ways to dynamically expand the computational resources of a transformer at inference time—increase the expressive power of transformers. Padding expands the circuit width of transformers, allowing them to resolve logical formulas over more variables. As a consequence of this, polynomially padded transformers can recognize exactly $\mathsf{TC}^{0}$, which was previously known only as an upper bound (Merrill and Sabharwal, 2023a; Pfau et al., 2024). In contrast, looping increases the depth of transformers. Applying looping on top of padding, we extended our result to show that logd-depth, padded transformers recognize exactly $\mathsf{TC}^{d}$. This means that transformers with polynomial padding and polylogarithmic looping converge to recognizing $\mathsf{NC}$, the largest class of problems that can be solved with parallel computation. In contrast, transformers with CoT have greater expressive power under standard complexity conjectures (Merrill and Sabharwal, 2023b), but suffer from slow sequential decoding. Thus, while looping and padding are not as powerful as CoT, our results suggest they are quite effective ways to expand transformers’ expressive power while preserving parallelism.
Several interesting open questions remain from this work. On the theoretical side, it would be valuable to develop a more finegrained characterization of $\mathsf{AHAT}^{d}_{k}$, i.e., looped transformers where the padding is at most $O(n^{k})$ for some fixed $k$. On the empirical side, while looped and padded transformers have both been explored already to an extent, it would be interesting to see whether these approaches could be successfully integrated to improve the performance of transformers on hard reasoning tasks. Further developing these approaches could ultimately lead to inference-time compute methods based on padding and looping that increase transformers’ expressive power without sacrificing parallelism, providing a more efficient alternative to CoT for solving moderately parallelizable problems.
Acknowledgments
Thanks to Selim Jerad, Andy Yang, and Michael Cadhilac, and attendees at the Formal Languages and Neural Networks (FLaNN) seminar for discussions. This project was supported by the National Science Foundation (NSF) through award 1922658 and WM’s NSF Graduate Research Fellowship. WM was also supported by a Two Sigma PhD Fellowship as well as by the Allen Institute for AI.
Appendix A Omitted Proofs
See 1
Proof.
That unmasked padded transformers can be simulated by mixed-masked transformers of the same depth holds trivially, as the former constitute a special case of the latter. We will next describe how to convert an unmasked padded transformer to a masked one, on a per-head basis. The construction will leave the computation of masked heads unchanged, making the approach suitable for converting both unmasked and mixed-masked padded transformers to masked ones.
Let $E$ be an unmasked encoder in $\mathsf{uAHAT}^{d}_{k}$. Then $E$ has depth $\ell=O(\log^{d}n)$ and operates over $n+O(n^{k})$ tokens (including original input and padding tokens). By Lemma 1, $E$ can be simulated by a causally masked decoder $D$ with depth $\ell$ and with $\ell\cdot(n+O(n^{k}))$ new padding tokens (after the original $O(n^{k})$ padding tokens used by $E$). Thus, the total number of padding tokens $D$ uses is $p=O(n^{k})+\ell\cdot(n+O(n^{k}))$. Since $\ell=O(\log^{d}n)$, this simplifies to $p=O(\log^{d}(n)\cdot n^{k^{\prime}})$ where $k^{\prime}=\max\{k,1\}$. This is $O(n^{k^{\prime}})$ when $d=0$, and $O(n^{k^{\prime}+1})$ when $d\geq 1$. This finishes the proof of the first two parts. For the third part, observe that by definition, $\mathsf{uAHAT}^{d}_{*}=\bigcup_{k=0}^{\infty}\mathsf{uAHAT}^{d}_{k}$. From the above, this in turn is contained in $\bigcup_{k=0}^{\infty}\mathsf{AHAT}^{d}_{k^{\prime}+1}\subseteq\mathsf{AHAT}^{d}_{*}$, as desired.
For the mixed-masking case, the construction in Lemma 1 will preserve the original computation of causally masked heads if we add an additional term with large negative weight $C$ (fixed w.r.t. $n$) that is activated if the index within the block is greater than $b_{i}$. We set $C$ large enough to dominate all other terms in the inner product computation at each token. As a result, the head is constrained to only attend to previous tokens within the block. Applying this modified construction on a case-by-case basis per head, we can take Proposition 1 to apply equally to transformers with mixed-masking. ∎
See 4
Proof.
As noted earlier, it is known that $\mathsf{AHAT}^{0}\subseteq\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}$ [Merrill and Sabharwal, 2023a, c, Chiang, 2025], and this generalizes to transformers with padding tokens [Pfau et al., 2024]. We will extend this construction to the case of $\mathsf{AHAT}^{d}_{*}$. The main observation is that even though the depth of a $O(\log^{d}n)$-depth padded AHAT transformer is input-dependent, the transformer has a looped architecture and thus a fixed (independent of input size) set of parameters. This makes it possible for a Turing machine (which can only store a fixed number of parameters) to construct an equivalent $\mathsf{TC}^{d}$ circuit family. [^3] The proof naturally generalizes to the case of logd-depth transformers that don’t necessarily have identical parameters across blocks of layers but whose parameters in each layer can be computed in an $\mathsf{L}$-uniform manner.
Let $L$ be a language in $\mathsf{AHAT}^{0}_{*}$ and let $T$ be a looped, padded AHAT transformer that, when unrolled $c\log^{d}n$ times for large enough $n$, [^4] As before, for small $n$, we assume $T$ uses a lookup table. recognizes whether $x\in\mathsf{L}$ for any input $x$ with $\lvert x\rvert=n$. Let $\langle A,B,C\rangle$ be the partition of layers of $T$ where $A$ is the set of initial layers, $B$ is the block that’s repeated $c\log^{d}n$ times on inputs of length $n$, and $C$ is the set of final layers. Each of these itself is a fixed-depth padded transformer; let’s call these $T_{A}$, $T_{B}$, and $T_{C}$. By prior results, there are $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit families $\{C^{A}_{n}\}_{n=0}^{\infty},\{C^{B}_{n}\}_{n=0}^{\infty},$ and $\{C^{B}_{n}\}_{n=0}^{\infty}$ that simulate transformers $T_{A},T_{B},$ and $T_{C}$, respectively. Thus, there is a (single) log-space Turing machine $M$ that, on input $\lvert x\rvert$, can construct each of these three circuit families. By looping the $T_{B}$ block $c\log^{d}n$ times, it can therefore also construct a circuit family $\{C_{n}\}_{n=0}^{\infty}$ that recognizes $L$. It follows that $L$ is in $\mathsf{L}$-uniform $\mathsf{TC}^{d}$. ∎
This argument does not work for $\mathsf{FO}$-uniform $\mathsf{TC}^{d}$ since it is not clearly that case that the circuit that repeats a layer $\log n$ times can be defined in $\mathsf{FO}$ just because the one that simulates one layer can [Merrill and Sabharwal, 2023c].
Appendix B $\mathsf{TC}^{d}$ Circuit Evaluation Problem
[…]
$\mathrm{Circuit}$ $\to\ \mathrm{Gate}^{*}$
$\mathrm{Gate}$ $\to\ \texttt{X}\;\mid\;\mathrm{Op}\;\mathrm{Arg}^{*}$
[…]
$\mathrm{Arg}$$\to\ \texttt{\&}\texttt{1}^{*}$
Semantically, we take the $k$-th X gate to return the $k$-th input bit. Other gates retrieve the values of the gates referred to by their argument pointers and apply the associated logical function. We take the final gate in the serialization to be the output gate of the circuit. Note that not all strings in this grammar represent a well-formed circuit, but any valid circuit can be serialized in this format.
As an example, the threshold circuit $\text{Majority}(x_{1},x_{2}\lor x_{3},\neg x_{3})$ stating that at least two of $x_{1}$, $x_{2}\lor x_{3}$, and $\neg x_{3}$ should be true, would be represented as follows:
$\underbrace{\texttt{X}\;\texttt{X}\;\texttt{X}}_{\text{input}}\;\underbrace{\texttt{MAJ}\;\texttt{\&1}\;\texttt{\&11111}\;\texttt{\&111111}}_{\text{Majority gate}}\;\underbrace{\texttt{OR}\;\texttt{\&11}\;\texttt{\&111}}_{\text{Or gate}}\;\underbrace{\texttt{NOT}\;\texttt{\&111}}_{\text{Not gate}}$
Note that the non-input gates need not be serialized in this particular order. The following is also an equally valid serialization of the same formula:
$\underbrace{\texttt{X}\;\texttt{X}\;\texttt{X}}_{\text{input}}\;\underbrace{\texttt{OR}\;\texttt{\&11}\;\texttt{\&111}}_{\text{Or gate}}\;\underbrace{\texttt{NOT}\;\texttt{\&111}}_{\text{Not gate}}\;\underbrace{\texttt{MAJ}\;\texttt{\&1}\;\texttt{\&1111}\;\texttt{\&11111}}_{\text{Majority gate}}$
We now formalize the circuit evaluation problem, for any class of circuit families, such as the class $\mathsf{TC}^{d}$ of logd-depth circuit families:
Definition 11 ($\mathsf{C}$ circuit evaluation).
Let $\mathsf{C}$ be a class of (potentially non-uniform) circuit families. The $\mathsf{C}$ circuit evaluation problem is defined as follows:
  • Input: $(x,\langle C\rangle)$ where $x\in\{0,1\}^{*}$ is a string and $\langle C\rangle$ is the serialization of a circuit $C$ such that $C=C_{\lvert x\rvert}$ for some circuit family $\{C_{n}\}_{n=0}^{\infty}\in\mathsf{C}$.
  • Output: The value $C(x)$.
We refer to the special case where $\mathsf{C}=\mathsf{TC}^{d}$ as the $\mathsf{TC}^{d}$ circuit evaluation problem. This corresponds to threshold circuit evaluation where the depth of the circuit is constrained to be $O(\log^{d}\lvert x\rvert)$.
We further define wide-$\mathsf{TC}^{d}\subseteq\mathsf{TC}^{d}$ as the class of circuit families $\{C_{n}\}_{n=0}^{\infty}$ such that there exists some $c$ such that, for large $n$, the depth of $C_{n}$ is at most $c\log^{d}n$ and, crucially, the size is at least $n^{c}$. That is, the size (and hence the width) of the circuit is large relative to its depth.
We define the corresponding wide-$\mathsf{TC}^{d}$ circuit evaluation problem as a variant of circuit evaluation problem with $\mathsf{C}=\text{wide-}\mathsf{TC}^{d}$.
B.1 Hardness of $\mathsf{TC}^{d}$ Circuit Evaluation
Lemma 5.
For $d\geq 1$, $\mathsf{TC}^{d}$ circuit evaluation is hard for $\mathsf{L}$-uniform $\mathsf{TC}^{d}$ under $\mathsf{L}$-reductions.
Proof.
Given any $L\in\mathsf{L}\text{-uniform}\ \mathsf{TC}^{d}$, there exists a log-space Turing machine $T_{L}$ that constructs a circuit family $\{C_{n}\}_{n=0}^{\infty}$ that recognizes $L$. We can construct an $\mathsf{L}$-reduction from $L$ to the $\mathsf{TC}^{d}$ circuit evaluation problem as follows. Given an input $x$ whose membership in $L$ we would like to check, the reduction first copies $x$ to the output. It then uses the log-space Turing machine $T_{L}$ to build the circuit $C_{\lvert x\rvert}$ and output it in the above serialized format. We thus have a log-space reduction from $x$ to $(x,\langle C_{\lvert x\rvert}\rangle)$. We conclude that $\mathsf{TC}^{d}$ circuit evaluation is hard for $\mathsf{L}$-uniform $\mathsf{TC}^{d}$ under $\mathsf{L}$-reductions. ∎
Lemma 6.
For $d\geq 1$, wide-$\mathsf{TC}^{d}$ circuit evaluation is hard for $\mathsf{L}$-uniform $\mathsf{TC}^{d}$ under $\mathsf{L}$-reductions.
Proof.
As in the proof of Lemma 5, we are given $L$ such that there exists a log-space Turing machine $T_{L}$ that constructs $\{C_{n}\}_{n=0}^{\infty}$ recognizing $L$, where the depth of each $C_{n}$ is at most $c\log^{d}n$ for some $c$ and large enough $n$. We can create a modified log-space Turing machine $T^{\prime}_{L}$ that builds $\{C^{\prime}_{n}\}_{n=0}^{\infty}$ that still recognizes $L$ in depth $c\log^{d}n$ but is padded with dummy gates so that it has size at least $n^{c}$: we do this by keeping a counter for the number of gates and wires produced and outputting dummy ones until $n^{c}$ is exceeded. We then follow the rest of the proof of Lemma 5 with $T^{\prime}_{L}$ instead of $T_{L}$. ∎
B.2 Solving $\mathsf{TC}^{d}$ Circuit Evaluation with Transformers
Lemma 7.
There is a mixed-masked looped transformer $T$ that, on input $(x,\langle C\rangle)$ where $x\in\{0,1\}^{*}$ and $\langle C\rangle$ is the serialization of a depth $\ell$ circuit with $\lvert x\rvert$ inputs, computes $C(x)$ when unrolled $\ell$ times.
Proof.
We adapt the proof of Merrill and Sabharwal [2023a, Theorem 3], which sketches how log-depth (unlooped) transformers can implement the $\mathsf{TC}^{0}$ circuit evaluation problem. We construct a looped transformer that will “attempt” to evaluate every gate: if its arguments have already been computed, the gate will return $\{0,1\}$, and, if not, it will return undefined ($\bot$).
[…]
X Gates. In the setup layers at an X token $i$, we use a causally masked uniform attention head with value $\mathbbm{1}[w_{j}=\texttt{X}]$. Thus, this head computes $r_{i}/i$, where $r_{i}$ number of X gates before and including $i$. We compute $\phi(r_{i}/i,1/i)=\phi(r_{i})$ and store it in the residual stream.
In the looped layers, we define an attention head with query $\phi(r_{i})$, key $\phi(j)$, and value $w_{j}$. This head thus retrieves input token $w_{r_{j}}$. We update the gate value to $v_{i}\leftarrow w_{r_{j}}$ (viewing both as vectors in the same space).
Other Gates. In the setup layers, each argument token $i$ attends with causally masked uniform attention with value $\mathbbm{1}[w_{j}=\&]$ to compute $a_{i}$, the number of arguments to its left (including it). Each & token attends with query $\phi(a_{i})$, key $\phi(a_{j})$, and value $\mathbbm{1}[w_{j}=\&]$, which returns $1/(1+z_{i})$, where $z_{i}$ is the number of 1’s following & token $i$. We compute and store $\phi(z_{i}+1)$ in the residual stream. We compute $g_{i}$ similarly to $a_{i}$, the number of gate tokens to left of token $i$ (also inclusive).
In the first looped layer, each & token $i$ attends with query $\phi(z_{i}+1)$, key $\phi(g_{j})$, and value $v_{j}$. Thus, the argument token $i$ retrieves $v_{z_{i}+1}$, the value at gate $z_{i}+1$. In the second looped layer, gate token $i$ attends with query $\phi(g_{i})$, key $\phi(g_{j})$, and value $v_{z_{j}+1}$. This, it returns the vector $\frac{1}{\lvert A_{i}\rvert}\sum_{j\in A_{i}}v_{j}$, where $A_{i}$ is the set of gates that are arguments of gate $i$. From this vector, we can apply projections to recover $T$, the fraction of $j\in A_{j}$ with $v_{j}=1$, as well as $U$, the fraction of $j\in A_{i}$ with $v_{j}=\bot$. If $U>0$, we set $v_{i}\leftarrow\bot$. Otherwise, we set $v_{i}$ by thresholding $T$ against some threshold $k$ based on the gate type ($T\geq 1$ for AND, $T>0$ for OR, and $T\geq 1/2$ for MAJ). In effect, this sets $v_{i}\leftarrow G(\{v_{j}\}_{j\in A_{j}})$, where $G$ is the gate type.
Thus, the looped layers either keep $v_{i}$ as $\bot$ or correctly update it to its true value. Furthermore, the number of looping steps until $v_{i}$ is updated is exactly the depth of node $i$. Thus, a circuit $C$ of depth $\ell$ can be fully evaluated by looping the depth-2 block $\ell$ times. ∎
Corollary 7.1.
For any $d\geq 0$, the wide-$\mathsf{TC}^{d}$ circuit evaluation problem is in $\mathsf{mAHAT}^{d}_{0}$, and hence in $\mathsf{AHAT}^{d}_{1}$.
Proof.
We are given input $(x,\langle C_{n}\rangle)$, where $C_{n}$ comes from some $\mathsf{TC}^{d}$ circuit family $\{C_{n}\}_{n=0}^{\infty}$ with depth at most $c\log^{d}n$ and size at least $n^{c}$, where $c$ is a constant specific to the circuit family. If we unroll the transformer from Lemma 7 to depth $c\log^{d}n$, we can solve wide-$\mathsf{TC}^{d}$ circuit evaluation problem for large enough $n$ (w.l.o.g. we can solve small-$n$ examples via table lookup).
We next justify that a mixed-masked mAHATd transformer will unroll at least $c\log^{d}n$ times for large enough $n$. This transformer will unroll exactly $\log^{d}N$ times, where $N=n+\lvert\langle C_{n}\rangle\rvert$ is the total input length for a circuit evaluation instance. Since the size of $C_{n}$ is at least $n^{c}$, we have that $N\geq n^{c}$. Thus, our $\mathsf{mAHAT}^{d}_{0}$ transformer unrolls the following number of times:
$\log^{d}N\geq\log^{d}n^{c}=c^{d}\log^{d}n\geq c\log^{d}n.$
Thus, our transformer will unroll a sufficient number of times to solve the wide-$\mathsf{TC}^{d}$ circuit evaluation problem. It follows that this problem is in $\mathsf{mAHAT}^{d}_{0}$.
Finally, we conclude via Proposition 1 that this mixed-masked transformer can be converted to a corresponding causally masked padded transformer, placing the problem also in $\mathsf{AHAT}^{d}_{1}$. ∎
Appendix C Uniformity Ceiling: Sanity Checking Our Results
We recount some simple but counterintuitive facts obtained while verifying the consistency of our results. In general, for classes $\mathsf{A},\mathsf{B}$ such that $\mathsf{A}\subseteq\mathsf{B}$, $\mathsf{B}$-uniformity often leads to larger classes of languages than $\mathsf{A}$-uniformity, as we have more resources to construct a circuit. However, as the following proposition shows, this is not always the case.
For the purposes of this result, we assume that $\mathsf{A}$-uniform $\mathsf{XC}$ means there exists $f\in\mathsf{A}$ such that $f$ maps $1^{n}$ to an $\mathsf{XC}$ circuit that operates over inputs of size $n$. This is not satisfied by some stronger notions of uniformity like $\mathsf{DLOGTIME}$ uniformity, but it is satisfied for $\mathsf{L}$ and $\mathsf{NL}$ uniformity.
Proposition 2 (Uniformity ceiling).
Consider classes $\mathsf{A},\mathsf{B}$ and a circuit class $\mathsf{XC}$ such that:
  1. 1.
    $\mathsf{A}\subseteq\mathsf{B}\subseteq\mathsf{A}\textrm{-uniform}\;\mathsf{XC}$;
  2. 2.
    $\mathsf{XC}$ circuit evaluation is in $\mathsf{A}$-uniform $\mathsf{XC}$.
Then $\mathsf{B}$-uniformity does not strengthen $\mathsf{A}$-uniformity, i.e., $\mathsf{A}\textrm{-uniform}\;\mathsf{XC}=\mathsf{B}\textrm{-uniform}\;\mathsf{XC}$.
Proof.
We will show that any language $L\in\mathsf{B}\textrm{-uniform}\;\mathsf{XC}$ is also in $\mathsf{A}\textrm{-uniform}\;\mathsf{XC}$. By definition of $\mathsf{B}\textrm{-uniform}\;\mathsf{XC}$, there exists a procedure $f\in\mathsf{B}$ that maps any input $\texttt{1}^{n}$ to an $\mathsf{XC}$ circuit $C^{f}_{n}$ that checks membership in $L$ over all strings $w$ of size $n$, i.e., $f(\texttt{1}^{n})=C^{f}_{n}$. By the first precondition of the proposition, $f$ is also in $\mathsf{A}\textrm{-uniform}\;\mathsf{XC}$. This means there exists a procedure $g\in\mathsf{A}$ that maps any input $\texttt{1}^{n}$ to an $\mathsf{XC}$ circuit $C^{g}_{n}$ that does what $f$ does, i.e., on input $\texttt{1}^{n}$ constructs the $\mathsf{XC}$ circuit $C^{f}_{n}$. Thus, we have $g(\texttt{1}^{n})=C^{g}_{n}$ and $C^{g}_{n}(\texttt{1}^{n})=C^{f}_{n}$. Furthermore, by the second precondition of the proposition, there exists a procedure $h\in\mathsf{A}$ that maps any input $(\texttt{1}^{n},\texttt{1}^{m})$ to an $\mathsf{XC}$ circuit $C^{h}_{n,m}$ that evaluates any $\mathsf{XC}$ circuit $C_{n}$ of size $m$ over any input $w$ of size $n$, i.e., $h(\texttt{1}^{n},\texttt{1}^{m})=C^{h}_{n,m}$ and $C^{h}_{n,m}(C_{n},w)=C_{n}(w)$.
Now we are ready to construct an $\mathsf{A}$-uniform $\mathsf{XC}$ circuit for recognizing $L$. Using currying notation from functional programming, [^5] Currying is the technique of translating a function that takes multiple arguments into a sequence of families of functions, each taking a single argument, cf. https://en.wikipedia.org/wiki/Currying. As an example, currying turns a two-argument function $f:(x,y)\mapsto z$ into a family of single-argument functions $g:x\mapsto(y\mapsto z)$, i.e., $g(x)$ is a single-argument function that maps $y$ to $z$ (in a manner that depends on $x$). this circuit is $C_{L}=h(\texttt{1}^{n},\texttt{1}^{m})(g(\texttt{1}^{n})(\texttt{1}^{n}))$ where $n=\lvert w\rvert$ and $m=\lvert g(\texttt{1}^{n})\rvert$. Here we use currying notation to specify that the first input of the two-input circuit $C^{h}_{n,m}$ generated by $h$ is fixed as $g(\texttt{1}^{n})(\texttt{1}^{n})$, turning it into a single-input function. Similarly, $g(\texttt{1}^{n})(\texttt{1}^{n})$ denotes first applying $g$ to $\texttt{1}^{n}$ in order to construct a circuit $g(\texttt{1}^{n})=C^{g}_{n}$, and then passing $\texttt{1}^{n}$ as input to this constructed circuit $C^{g}_{n}$.
By design, $C_{L}$ takes as input a string $w$ and computes $C_{L}(w)=h(\texttt{1}^{n},\texttt{1}^{m})(g(\texttt{1}^{n})(\texttt{1}^{n}),w)$.
It is easy to see that $C_{L}$ can be constructed by a procedure in $\mathsf{A}$—after all, $g$ and $h$ are in $\mathsf{A}$, $\texttt{1}^{n}$ can be easily computed from $w$ by converting all 0s to 1s, and similarly $\texttt{1}^{m}$ can be computed from the output $g(\texttt{1}^{n})$ of applying procedure $g$ (which is in $\mathsf{A}$) on $\texttt{1}^{n}$. Further, $C_{L}$ is a circuit in $\mathsf{XC}$ as it is a fixed composition of the $\mathsf{XC}$ circuits that functions $h$ and $g$ output. All that remains to be shown is that $C_{L}$ recognizes $L$. To this end, we note:
$C_{L}(w)$ $=h(\texttt{1}^{n},\texttt{1}^{m})(g(\texttt{1}^{n})(\texttt{1}^{n}),w)$
$=C^{h}_{n,m}(g(\texttt{1}^{n})(\texttt{1}^{n}),w)$
$=C^{h}_{n,m}(C^{g}_{n}(\texttt{1}^{n}),w)$
$=C^{h}_{n,m}(C^{f}_{n},w)$
$=C^{f}_{n}(w).$
By design, $C_{f}(w)=1$ if and only if $w\in L$. Thus, $C_{L}$ recognizes $L$. ∎
Applying this result in the case of $\mathsf{L}$ and $\mathsf{NL}$ as classes $\mathsf{A}$ and $\mathsf{B}$, respectively, we obtain:
Corollary 3.1.
For any circuit class $\mathsf{XC}$ containing $\mathsf{NL}$, $\mathsf{L}\textrm{-uniform}\;\mathsf{XC}=\mathsf{NL}\textrm{-uniform}\;\mathsf{XC}$.
This may appear surprising on the surface level, but it has a quite intuitive interpretation: weaker uniformity provably cannot increase a circuit class that is sufficiently expressive. For example, applying Proposition 2 with $\mathsf{A}=\mathsf{L}$, $\mathsf{B}=\mathsf{P}$, and $\mathsf{XC}=\mathsf{P}$ (poly-size circuits), we can derive the unsurprising result that relaxing the uniformity condition for polynomial-size circuits does not increase their expressive power beyond $\mathsf{P}$:
Corollary 3.2.
$\mathsf{L}\textrm{-uniform}\;\mathsf{P}=\mathsf{P}\textrm{-uniform}\;\mathsf{P}$